OpenStudy (princesssleelee):

Solve the following system of equations. 5x-7y=17 −7x+9y=−19 x=___ y=___

OpenStudy (anonymous):

we can do it, it is going to take a couple steps and the answer is messy if you want the y terms to drop out, multiply the first equation all the way across by \(9\) and the second all they way across by \(7\) let me know what you get

OpenStudy (princesssleelee):

what?

OpenStudy (anonymous):

multiply each term in \[2x-7y=17\] by \(9\)

OpenStudy (anonymous):

oops i meant in \[5x-7y=17\]

OpenStudy (princesssleelee):

18x-63y=153

OpenStudy (princesssleelee):

45x

OpenStudy (anonymous):

right

OpenStudy (princesssleelee):

-49+81=-133

OpenStudy (anonymous):

\[45x-63y=153\]

OpenStudy (anonymous):

now multiply each term in the next equation by \(7\)

OpenStudy (princesssleelee):

i did

OpenStudy (princesssleelee):

-49+81=-133

OpenStudy (anonymous):

\[7\times 9=63\]

OpenStudy (anonymous):

not 81

OpenStudy (princesssleelee):

sorry i was thinking 9x9 still lol

OpenStudy (anonymous):

i know

OpenStudy (anonymous):

u should get \[-49x+63y=-133\]

OpenStudy (princesssleelee):

correct

OpenStudy (anonymous):

you have now \[45x-63y=153\\ -49x+63y=-133\] now add in a column and the y terms will go bye bye

OpenStudy (princesssleelee):

what lol

OpenStudy (anonymous):

add the first line and the second line

OpenStudy (anonymous):

what is \(45x-49x\)?

OpenStudy (princesssleelee):

x=-10/47?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

did you see the link i sent?

OpenStudy (princesssleelee):

lol yes are those correct?

OpenStudy (anonymous):

yes technology is wonderful

OpenStudy (anonymous):

you can see what i wrote in the link, just copied and pasted what you wrote, but put a comma between them

OpenStudy (princesssleelee):

ok no they are not you copied and pasted the last question that i wrote incorrectly, the 5x is not negative. its x=-5 y=-6

OpenStudy (anonymous):

i believe you